OpenStudy (anonymous):

Given f(x) = 2x4- 3x2 - 8x + 4 , what is f(3)? A. 88 B. 115 C. 131 D. 133

OpenStudy (michele_laino):

here, you have to replace \(x=3\), so, we have: \[\Large f\left( 3 \right) = 2 \cdot \left( {{3^4}} \right) - 3 \cdot \left( {{3^2}} \right) - 8 \cdot 3 + 4 = ...?\]

OpenStudy (anonymous):

Yeah. So do I just solve all of that?

OpenStudy (michele_laino):

yes! please keep in mind that: \(3^4=81\), \(3^2=9\)

OpenStudy (anonymous):

...I got 130.

OpenStudy (anonymous):

Oh crap nevermind hold on.

OpenStudy (anonymous):

115?

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

Thank you for your help!

OpenStudy (michele_laino):

:)

OpenStudy (anonymous):

Do you think you could help me on my other problems?

OpenStudy (anonymous):

@Michele_Laino